OpenStudy (anonymous):

the difference of two numbers is 1 and the sum of their squares is 41, find the numbers

OpenStudy (anonymous):

You have to translate the words into math symbols.

OpenStudy (anonymous):

ok i know that x-x=1

OpenStudy (anonymous):

and the exponents have to give me 41

OpenStudy (anonymous):

For your first response, it would be better if you use a different variable for each number rather than x and x.

OpenStudy (anonymous):

x-y=1

OpenStudy (anonymous):

You are pretty good at this. You need to do the same thing for the second part.

OpenStudy (anonymous):

ok but

OpenStudy (anonymous):

Yah

OpenStudy (anonymous):

what would the answer be

OpenStudy (tamtoan):

if your variables are x and y, how do you write the sum of their square ?

OpenStudy (anonymous):

aaahh i dont know

OpenStudy (anonymous):

(5,4)?

OpenStudy (tamtoan):

if the variable is x, the square of that variable is \[x^{2}\] likewise with the other variable ...and the sum of both is ? ...try to write that expression

OpenStudy (anonymous):

x^2-y^2=1

OpenStudy (anonymous):

2+2?

OpenStudy (anonymous):

no no

OpenStudy (tamtoan):

the sum, not the difference ...and the sum of them is 41..do that again ?

OpenStudy (anonymous):

4 and 5

OpenStudy (tamtoan):

where do you get 4 and 5 ?

OpenStudy (anonymous):

five minus four is 1 and 4^2+5^2 is 41

OpenStudy (tamtoan):

that's guessing, you guess right, but it's not the way to solve :) you have to write a system of 2 equations and try to solve it.